Find the slope and the y-intercept of the line
10x=3y+6
Thank you

Answer by haileytucki(390)

10x=3y+6
Since y is on the right-hand side of the equation, switch the sides so it is on the left-hand side of the equation.
3y+6=10x
Since 6 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 6 from both sides.
3y=-6+10x
Move all terms not containing y to the right-hand side of the equation.
3y=10x-6
Divide each term in the equation by 3.
(3y)/(3)=(10x)/(3)-(6)/(3)
Simplify the left-hand side of the equation by canceling the common factors.
y=(10x)/(3)-(6)/(3)
Simplify the right-hand side of the equation by simplifying each term.
y=(10x)/(3)-2
To find the slope and y intercept, use the y=mx+b formula where m=slope and b is the y intercept.
y=mx+b
Using the y=mx+b formula, m=(10)/(3) and b=-2.
m=(10)/(3), b=-2